OpenStudy (anonymous):

solve: 7y+4=3-2y.. having a brain freeze

OpenStudy (lgbasallote):

solve for y?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

7y + 4 = 3 - 2y first add 2y to both sides 9y + 4 = 3 now subtract 4 from both sides 9y = -1 now divide both sides by 9 y = -1/9 does that help?

OpenStudy (anonymous):

http://www.myalgebra.com/algebra_solver.aspx use this and plug in your equation. Its a lifesaver for checking your answer too

OpenStudy (anonymous):

yes i completely skipped the divide by 9 step... whoops. Thanks(:

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

i will check it out... thanks!!

OpenStudy (anonymous):

@candrie ...amazing website!! will use to check my work thanks!

OpenStudy (anonymous):

yes I know!! Its a total lifesaver! My friend told me about it